OpenStudy (anonymous):

Need a refresher here 6u^2-6/u^2-6u-7

OpenStudy (rajee_sam):

\[\frac{ 6u ^{2} - 6}{u ^{2} -6u - 7 }\] Factor the top and bottom and cancel out the common factors and simplify

OpenStudy (rajee_sam):

top can be factored as \[6 (u ^{2} - 1) = 6 (u+1) (u - 1)\] Difference of Squares. Factor the denominator like regular Quadratic function

OpenStudy (rajee_sam):

\[u ^{2} - 6u - 7 = (u-7) (u + 1)\]

OpenStudy (rajee_sam):

now cancel out the common factors and write the rest as your answer.

OpenStudy (rajee_sam):

\[\frac{ 6(u + 1) (u - 1) }{ (u - 7 ) ( u + 1) }\]

OpenStudy (anonymous):

so would it be 6(u-1)/(u-7) ?

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

Sweet thank you.